OpenStudy (anonymous):

OK, I need someone who can show me step by step how to solve this equation (put it to this form ax2+bx+c) : Kc = ((2x)^2 * (2x)^2) /((0.53-2x)^2 * (0.515-x)) = 9.14 * 10^-6 Thank you very much! cant get the right answere!

OpenStudy (aaronq):

it's just math \(\dfrac{(2x)^2 * (2x)^2}{ (0.53-2x)^2 * (0.515-x)} = 9.14 * 10^{-6}\) \(\dfrac{4x^2 * 4x^2}{ (0.53-2x)*(0.53-2x) * (0.515-x)} = 9.14 * 10^{-6}\) \(\dfrac{16x^4}{ (0.2809-2.12x +4x^2)* (0.515-x)} = 9.14 * 10^{-6}\) etc

OpenStudy (aaronq):

keep expanding the bottom first

OpenStudy (anonymous):

Thank you! But now I got this: 16x^4/(0.145 - 1.373x + 4.18x^2 - 4x^3) what do I do next?

OpenStudy (aaronq):

are you sure you can't omit some of the values to simplify it? you're ending up with a quartic function thats pretty much impossible to solve by hand lol